### v2

# A theory of phonon-induced friction on molecular adsorbates

Abstract
In this manuscript, we provide a general theory for how surface phonons couple to molecular adsorbates. Our theory maps the extended dynamics of a solid’s atomic vibrational motions to a generalized Langevin equation, and by doing so captures these dynamics in a single quantity: the non-Markovian friction. The different frequency components of this friction are the phonon modes of the solid weighted by their coupling to the adsorbate degrees of freedom. Using this formalism, we demonstrate that physisorbed species couple primarily to acoustic phonons while chemisorbed species couple to dispersionless local vibrations. We subsequently derive equations for phonon-adjusted reaction rates using transition state theory and demonstrate that these corrections improve agreement with experimental results for CO desorption rates from Pt(111).
1 Introduction
The atomic vibrations of a solid surface can significantly influence the rates and mechanisms of surface chemical processes.1, 2, 3, 4 Unraveling these influences experimentally is difficult, necessitating the development of theoretical tools. In this manuscript, we introduce a theory for the coupling between the nuclear dynamics of surface-bound (adsorbed) molecules and the vibrations of the underlying surface. Our theory projects the collective surface vibrations onto the motion of the adsorbate, and in doing so captures how surface phonons spanning several order of magnitude in length/energy scale influence adsorbate dynamics.
Over the past four decades, a multitude of experimental and theoretical studies have demonstrated how surface vibrations affect reaction dynamics at solid interfaces. For example, several studies have shown that surface atom motion plays a central role in the dissociation of methane and molecular nitrogen on a variety of different metal surfaces. 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17. Certain experiments have utilized piezoelectric drivers to generate surface acoustic waves and demonstrated that waves of particular frequencies and polarizations are capable of enhancing some reactions, such as the oxidation of ethanol or carbon monoxide.18, 19, 20, 21, 22, 23, 24, 25. While several mechanisms have been proposed to explain this rate enhancement, the origin of this effect remains largely uncertain.25 Much of this uncertainty stems from the inability of current theory to connect the lengthscales and timescales of atomistic dynamics with the mesoscopic scale of experimentally relevant surface acoustic waves 26. The theoretical approach that we present here is capable of effectively connecting these scales and thereby providing new physical insight into the roles of surface vibrations in surface chemical reactions.
Our theoretical model describes the motion of surface-bound molecules via a generalized Langevin equation (GLE). The central parameter of this equation is the non-Markovian friction kernel (or memory kernel), which encapsulates the different surface vibrational modes weighted by their coupling to the adsorbate. Using this formalism, we demonstrate that the influence of surface phonons on the dynamics of an adsorbed molecule depends significantly on the magnitude of the adsorbate-surface coupling. Specifically, we find that chemisorbed species couple primarily to dispersionless local vibrations, while physisorbed species couple primarily to acoustic phonons across a broad range of frequencies. The key parameter that determines whether the primary influence of surface vibrations is via extended or localized phonon modes is the ratio of the solid’s Debye frequency to the frequency of the adsorbate-surface bond (Figure 1). By combining these observations with harmonic transition state theory, we derive equations that describe how phonons alter reaction rates at solid surfaces and demonstrate that these phononic corrections agree with experimental measurements of desorption rate constants.
Figure 1: 2D schematic illustrating the dominant phonon modes in terms of their coupling to the adsorbate or contribution to the friction kernel. $\omega_{\mathrm{as}}$ is the frequency of the adsorbate-surface bond and the $\omega_{\mathrm{D}}$ is the solid’s Debye frequency.
The approach we present in this manuscript extends our previous method for utilizing the GLE for describing the influence of surface vibrational modes on molecular adsorbates. farahvash_using_2023 In our previous method, the GLE was designed to describe the collective influences of the surrounding solid on a single representative surface atom (i.e., binding site). In the method we present here, the GLE is designed to describe the influence of the entire surface on the dynamics of the adsorbate itself. This new approach is more easily interpretable because the friction kernel directly represents influence of surface phonons on the adsorbate.
The remainder of the paper is organized as follows. In Section 2, we present the formal theory behind the phonon-induced GLE and highlight the characteristics of the friction kernel. Along with the derivation of our theoretical method, we present its application to a simple model system of an adsorbate whose surface bond has a tunable frequency. We then analyze results for the friction kernel across a range of values for adsorbate-surface bond frequency. In Section 3, we provide further insight into the results presented in Section 2 by analyzing the role of dispersion plays in the phonon-induced friction. Finally, in Section 4, we develop a formulation of the transition state theory desorption rate that explicitly accounts for the influence of surface phonon modes. We validate this formulation by demonstrating an improved agreement with experiments of CO desorption from Pt(111).
2 Coupling surface vibrations to adsorbates
2.1 Theory
In this section, we derive a theory for modeling the effects of solid surface vibrations on the dynamics of an adsorbate. The general approach involves integrating out the nuclear degrees of freedom of the solid and representing their influence as effective equations of motion for the adsorbate. In order to do so we must make two critical assumptions. First, we assume that the nuclear degrees of freedom of the adsorbate and solid can be described classically and adiabatically. There are cases where adsorbate dynamics are non-adiabatic, and in such cases a mixed quantum-classical GLE approach has been shown to be effective 27, 28, however the development and analysis of such a theory is beyond the scope of this paper. Second, we assume that the solid potential energy surface and surface-adsorbate interaction are both harmonic. This approximation simplifies the formulation and is appropriate for the development of generalizable physical insight. We note that the effects of anharmonicity can be incorporated into the theory, which would be a natural target of future developments.
To begin, we define $\mathbf{x}_{\mathrm{A}}$ and $\mathbf{x}_{\mathrm{S}}$ as the the mass-weighted displacements of the adsorbate and solid nuclei, respectively, from their equilibrium positions. We decompose the total potential energy into contributions from the adsorbate and solid,
$V(\mathbf{x}_{\mathrm{A}},\mathbf{x}_{\mathrm{S}})=V_{\mathrm{A}}(\mathbf{x}_{\mathrm{A}})+V_{\mathrm{AS}}(\mathbf{x}_{\mathrm{A}},\mathbf{x}_{\mathrm{S}})+V_{S}(\mathbf{x}_{\mathrm{S}}),\$ (1)
where $V_{\mathrm{A}}$ is the adsorbate potential energy containing all intramolecular and intermolecular interactions, $V_{\mathrm{S}}$ is the solid potential energy surface, and $V_{\mathrm{AS}}$ the adsorbate-solid interaction. A second-order expansion of $V_{\mathrm{AS}}$ and $V_{\mathrm{S}}$ around the equilibrium nuclear positions invites the definition of the matrices $\mathbf{G}_{\mathrm{AS}}$, $\mathbf{G}_{\mathrm{A}}$, $\mathbf{G}_{\mathrm{S}}$, and $\mathbf{H}_{\mathrm{S}}$, whose elements are given by,
$G_{\mathrm{AS};ij}=\frac{\partial^{2}V_{\mathrm{AS}}}{\partial{x}_{\mathrm{A};i}\partial{x}_{\mathrm{S};j}},$ (2)
$G_{\mathrm{A};ij}=\frac{\partial^{2}V_{\mathrm{AS}}}{\partial{x}^{2}_{\mathrm{A};i}}\delta_{ij},$ (3)
$G_{\mathrm{S};ij}=\frac{\partial^{2}V_{\mathrm{AS}}}{\partial{x}^{2}_{\mathrm{S};i}}\delta_{ij},$ (4)
$H_{\mathrm{S};ij}=G_{\mathrm{S};ij}+\frac{\partial^{2}V_{\mathrm{S}}}{\partial{x}_{\mathrm{S};i}\partial{x}_{\mathrm{S};j}},$ (5)
where $\delta_{ij}$ is the Kronecker delta, $\mathbf{H}_{\mathrm{S}}$ is the solid’s mass-weighted Hessian, $\mathbf{G}_{\mathrm{AS}}$ is the coupling between adsorbate and solid degrees of freedom, and $\mathbf{G}_{\mathrm{A}}$ and $\mathbf{G}_{\mathrm{S}}$ are diagonal matrices of oscillation frequencies for the adsorbate and solid degrees of freedom respectively. All derivatives in Eqs. 2-5 are evaluated the equilibrium nuclear positions.
The expansion of $V_{\mathrm{AS}}$ and $V_{\mathrm{S}}$ can be manipulated to generate equations of motion for adsorbate,
$\accentset{\mbox{\large..}}{\mathbf{x}}_{\mathrm{A}}=-\frac{\partial V_{\mathrm{A}}}{\partial\mathbf{x}_{\mathrm{A}}}-\mathbf{G}_{\mathrm{A}}\mathbf{x}_{\mathrm{A}}-\mathbf{G}_{\mathrm{AS}}\mathbf{x}_{\mathrm{S}},\$ (6)
and solid atoms,
$\accentset{\mbox{\large..}}{\mathbf{x}}_{\mathrm{S}}=-\mathbf{G}^{T}_{\mathrm{AS}}\mathbf{x}_{A}-\mathbf{H}_{S}\mathbf{x}_{\mathrm{S}}.$ (7)
By utilizing a second-order expansion, we impose a harmonic approximation on the model. It is convenient to describe the nuclear displacements of the solid in a basis of phonon modes. Thus, we introduce solid phonon coordinates, $\mathbf{u}_{\mathrm{S}}=\mathbf{U}^{T}\mathbf{x}_{\mathrm{S}}$, here $\mathbf{U}$ is a matrix with the eigenvectors of $\mathbf{H}_{\mathrm{S}}$ as columns. In terms of these coordinates the equations of motion may be expressed as,
$\ddot{\mathbf{x}}_{\mathrm{A}}=-\frac{\partial V_{\mathrm{A}}}{\partial\mathbf{x}_{\mathrm{A}}}-\mathbf{G}_{\mathrm{A}}\mathbf{x}_{\mathrm{A}}-\mathbf{C}\mathbf{u}_{\mathrm{S}},$ (8)
$\ddot{\mathbf{u}}_{\mathrm{S}}=-\mathbf{C}^{T}\mathbf{x}_{\mathrm{A}}-\boldsymbol{\omega}^{2}\mathbf{u}_{\mathrm{S}},$ (9)
where $\mathbf{C}=\mathbf{G}_{\mathrm{AS}}\mathbf{U}$ is a matrix of couplings between each adsorbate coordinate and solid phonon mode, and $\boldsymbol{\omega}^{2}=\mathbf{U}^{T}\mathbf{H}_{\mathrm{S}}\mathbf{U}$ is a diagonal matrix of the squared frequencies of the solid phonons. Combining Eq. 8 and Eq. 9 yields a GLE for the adsorbate degrees of freedom29, 30,
[…]
where the friction kernel $\mathbf{K}(t)$ is given by,
$\mathbf{K}(t)=\mathbf{C}\frac{\cos(\boldsymbol{\omega}t)}{\boldsymbol{\omega}^{2}}\mathbf{C}^{T},$ (11)
and the statistics of the random force $\mathbf{R}(t)$ are related to $\mathbf{K}(t)$ by the second fluctuation-dissipation theorem,
$\frac{\left\langle\mathbf{R}(t)\mathbf{R}^{T}(0)\right\rangle}{k_{\mathrm{B}}T}=\mathbf{K}(t).$ (12)
The positive frequency components of the Fourier transform of the friction kernel are (up to a multiplicative constant),
$\bar{K}_{ab}(\omega)=\sum_{j}\frac{C_{aj}C_{bj}}{\omega^{2}_{j}}\delta{(\omega-\omega_{j})},$ (13)
where $\omega_{j}$ is the frequency of $j$th phonon mode of the solid. Alternatively, we can express this quantity as,
$\bar{K}_{ab}(\omega)=\frac{C_{a}(\omega)C_{b}(\omega)}{\omega^{2}}\rho(\omega),$ (14)
where $\rho(\omega)$ is the density of phonon modes at frequency $\omega$.
Some insight can be derived by studying the formulation above. Equation 10 illustrates that within the harmonic approximation, the influence of phonon modes on adsorbates can be described entirely in terms of $\mathbf{K}(t)$, as it determines both the properties of $\mathbf{R}(t)$ via the fluctuation-dissipation theorem and the deviation of the adsorbate potential from that of an adsorbate on a fixed solid. Equations 13 and 14 reveal that the Fourier transform of $\mathbf{K}(t)$ can be interpreted as the phonon density of states reweighted by the value of the adsorbate-phonon coupling $\mathbf{C}$. Indeed, $\mathbf{\bar{K}}_{ab}(\omega)$ is often referred to as the spectral density of the environment.
In some cases Eqs. 11 and 13 may further be simplified by stating that adsorbates couple only to a handful of surface sites and thus the matrices $\mathbf{G}_{AB}$ and $\mathbf{C}$ are sparse. For example, consider the case of an adsorbate that interacts with a single surface atom via a single adsorbate atom (e.g. CO on Pt(111) 31, 32). If the force constant associated with this interaction is $\mu\omega_{\mathrm{as}}^{2}$, where $\mu$ is the reduced mass of the adsorbate and surface atom, Eqs. 11 and 13 may be simplified to,
$K(t)=\frac{\mu^{2}}{mM}\omega_{\mathrm{as}}^{4}\sum_{j}\frac{U^{2}_{sj}}{\omega^{2}_{j}}\cos(\omega_{j}t),$ (15)
$\bar{K}(\omega)=\frac{\mu^{2}}{mM}\omega_{\mathrm{as}}^{4}\sum_{j}\frac{U^{2}_{sj}}{\omega^{2}_{j}}\delta{(\omega-\omega_{j})},$ (16)
where $U_{sj}$ is the expansion coefficient of the surface displacement $s$ in the $j$th normal mode, $m$ is the mass of the adsorbate atom, and $M$ is the mass of the surface atom.
The formulation above provides a framework for representing the influence of surface phonons on an adsorbate implicitly via the GLE. With a well-defined model of an adsorbate, surface, and their interactions, the friction kernel can be calculated. In the following subsection, we illustrate the application of this formalism to simple model systems.
2.2 Results and discussion
We have computed the phonon friction kernel for a simple model of CO on Pt(111). The CO was modeled as a single adatom and assumed the interact with a single adsorption site, as experimental structures show that CO adsorbs primarily atop Pt(111) sites 31, 32. We artificially vary the adsorbate-surface frequency $\omega_{\mathrm{as}}$ over a range of physically motivated values, in order asses how the strength of the adsorbate-surface interaction affects the properties of the friction kernel. The metal potential energy surface $V_{S}$ was modeled using an effective medium theory (EMT) forcefield developed by Norskov et al.33 More details about calculations can be found in the Section 5.
Figure 2: (A) Friction kernel, (B) spectral density, and (C) density of states for Pt(111) and three values of $\omega_{\mathrm{as}}$. The density of states in (C) was calculated using the 4x4x8 surface slab.
Figure 2 illustrates the results of our calculations for a selected set of $\omega_{\mathrm{as}}$ values ranging from those characteristic of weakly adsorbed species to those characteristic of strongly adsorbed species. We plot the friction kernel computed in the direction normal to the surface in systems with either a 4x4x8 or 8x8x8 Pt(111) slab. The lower values of $\omega_{\mathrm{as}}$ that we consider lie below that of the platinum’s Debye frequency of $\omega_{\mathrm{D}}=156\mathrm{cm}^{-1}$. We associate these values with the physisorbed regime, as they are characteristic of weakly adsorbed systems such as noble gases on Pt(111).34 The higher values of $\omega_{\mathrm{as}}$ that we consider are associated with the chemisorbed regime, corresponding to physical systems such as CO on Pt(111), which has $\omega_{\mathrm{as}}\approx 480\ \mathrm{cm}^{-1}$).32, 35, 36 Delta function peaks in $\bar{K}(\omega)$ and $\rho(\omega)$ were broadened to thin Lorentzians of width $1\mathrm{cm}^{-1}$ for ease of visualization.
For all cases of $\omega_{\mathrm{as}}$, the adsorbate couples strongly to a surface acoustic mode appearing around $\omega=10\mathrm{cm}^{-1}$. The location of this peak is highly dependent on the dimensions of the simulated surface slab, consistent with the behavior of an acoustic phonon. In Sec. 3, we give a more detailed theoretical and numerical analysis of the size effects observed here, fully accounting for phonon dispersion. In the physisorbed regime, the friction kernel is dominated by this acoustic mode, resulting in highly non-Markovian behavior. The phonon mode associated with this peak is associated with the flexing of the lattice in the direction perpendicular to the surface. Snapshots of this motion are provided in the insets in Figure 1 and video animations are provided in the electronic SI.
In the chemisorbed regime, the friction kernel is dominated by a high frequency mode around $\omega=210\ \mathrm{cm}^{-1}$. This mode arises from the local oscillations of the surface site the adsorbate is bound to. The frequency of this mode is independent of simulation cell size, suggesting that it is dispersionless. The large amplitude of this mode signifies that in the chemisorbed limit, the adsorbate is primarily sensitive to the local oscillations of the surface binding site (which are shifted in frequency due to the presence of the adsorbate). Many published models for reactive scattering on solid surfaces will describe solid vibrations using only a single, harmonically bound surface atom 6, 37, 38, 39, 40. These results explain why such a method is successful for strongly coupled species.
Despite this large shift in the spectral density when varying $\omega_{\mathrm{as}}$, in Figure 2C we demonstrate that the phonon density of states is nearly identical to that of a bare surface. The only significant change in $\rho(\omega)$ is the presence of the aforementioned surface site local mode present around $\omega=210\ \mathrm{cm}^{-1}$. These results make physical sense given that the low frequency acoustic modes of a solid should be unaffected by gaseous species, especially at low pressures and surface coverages.
To evaluate the effect of the solid force field and structure on these results, we have tested several different crystal structures, facets, elemental compositions, and force fields. The results of these tests are presented in the Section S1 of the SI. While the quantitative properties of the friction kernel vary across different systems (e.g. Ru has a much higher Debye frequency than Pt), the qualitative dependence of the friction kernel on $\omega_{\mathrm{as}}$ is quite general. This independence with respect to the details of the atomistic model is appealing and can be understood from a perturbative perspective. Specifically, in the physisorbed limit, the motion of the surface binding site is a small perturbation to the bulk phonon modes, while in the chemisorbed limit, the bulk phonon modes are a small perturbation to the motion of the surface binding sites. In the SI, we rigorously examine this statement by comparing perturbative schemes to the exact results for $K(t)$ and $\bar{K}(\omega)$.
Finally, we note that in this section we have only studied the effects of phonons for an ideal clean surface. Presumably, sources of surface heterogeneity, such as steps and defects, may add further richness and complexity to the picture we have provided here. We also note that we have only studied elemental solids, and the optical modes of polyatomic crystals could add another interesting dimension to this physical picture.
3 Phonon dispersion
The dependence of the acoustic peak frequency on simulation size, such as illustrated in Fig. 2, is an artifact that has the potential to prevent straightforward comparison between experiment and simulation. With or without periodic boundary conditions, the size of the simulated solid limits the maximum wavelength surface phonon. In this section, we illustrate how to generalize the theory presented in Section 2 by computing the friction kernel in the limit of an infinite surface through integration of frequencies and eigenmodes across the first Brillouin zone.
In an infinite crystalline solid, displacement by a lattice vector returns the same solid. This symmetry can be leveraged to calculate the phonon frequencies and displacements of the bulk solid using a spatial Fourier transform of the mass-weighted Hessian. This approach leads to the well-known Bloch’s theorem, which is summarized as follows. Let $a$ and $b$ be the indices of two primitive unit cells, and let $\mathbf{H}_{S}(a,b)$ be the mass-weighted Hessian of the crystal, ${H}_{S;ij}(a,b)=\frac{\partial^{2}V}{\partial{x}_{S,i}(a)\partial{x}_{S,j}(b)}$. The Fourier transform of this matrix may be expressed as,
$\mathbf{D}(\mathbf{k})=\sum_{a,b}\mathbf{H}_{\mathbf{S}}(a,b)e^{i\mathbf{k}\cdot(\mathbf{r}_{a}-\mathbf{r}_{b})},$ (17)
where $\mathbf{r}_{a}$ is the origin of the $a$th cell. $\mathbf{D}(\mathbf{k})$ is known as the dynamical matrix satisfying,
$\mathbf{D}(\mathbf{k})\mathbf{U}_{j}(\mathbf{k})=-\omega_{j}^{2}(\mathbf{k})\mathbf{U}_{j}(\mathbf{k}).$ (18)
where $\omega_{j}$ is the $j$th phonon band frequency and $\mathbf{U}_{j}$ is the corresponding polarization vector. For 3D monatomic crystals, the primitive (Wigner-Seitz) cell consists of a single atom with three degrees of freedom, thus producing three phonon bands. These bands, illustrated in Fig. 3 for FCC platinum, are the acoustic transverse and longitudinal modes, and all characterized by linear dispersion at low wavenumber.
In a surface slab, the presence of anisotropy breaks the symmetry of the 3D crystal. This symmetry breaking leads to additional surface mode bands, which can be acoustic (Rayleigh waves), or non-acoustic in character41, 42, 43. In Fig. 4B. and 4C. we demonstrate the phonon dispersion of a Pt(111) surface calculated via EMT. These results were calculated using a 4x4x8 surface replicated in a 6x6 super-cell. We verified convergence with respect to supercell size. Of course, the number of surface phonon bands depends on the size of the surface unit cell used in computing $\mathbf{D}(\mathbf{k})$, however, we demonstrate in in Figure S7 results are qualitatively similar across different unit cell sizes.
Figure 3: Phonon dispersion curves. (A) Bulk Pt dispersion curves calculated using an EMT forcefield and a 10x10x10 atom supercell. (B) Dispersion curves for a 4x4x8 atom surface slab replicated in a 6x6 surface cell (C) Same as middle but with an adsorbed CO molecule corresponding to $\omega_{\mathrm{as}}=480\ \mathrm{cm}^{-1}$.
Figure 3B illustrates the phonon dispersion for a bare surface, while Figure 3C illustrates the dispersion for a surface with an adsorbed CO. The bands are colored based on their mean slope between the $\Gamma$ and $K$ high-symmetry points. The three lowest frequency bands are all surface acoustic modes, as exemplified by the flexing mode depicted in the inset of Fig. 1, confirming that the corresponding peak in the phonon spectral density in Fig. 2 arises from an acoustic phonon. The remaining modes are nearly dispersionless — especially the highest frequency mode in Fig. 3C, which corresponds to the surface site local vibration. The dispersionless nature of this mode supports why it was not seen to be dependent on surface slab size in Fig. 2.
Using these dispersion relations, we can average the friction kernel across the first Brillouin zone,
$K(t)=\frac{\mu^{2}}{mM}\omega_{\mathrm{as}}^{4}\sum_{j}{\sum_{\mathbf{k}}}^{\prime}\ \frac{\left|U_{sj}(\mathbf{k})\right|^{2}}{\omega_{j}(\mathbf{k})^{2}}\cos(\mathbf{\omega}_{j}(\mathbf{k})t),$ (19)
$\bar{K}(\omega)=\frac{\mu^{2}}{mM}\omega_{\mathrm{as}}^{4}\sum_{j}{\sum_{\mathbf{k}}}^{\prime}\ \frac{\left|U_{sj}(\mathbf{k})\right|^{2}}{\omega_{j}(\mathbf{k})^{2}}\delta(\omega-\mathbf{\omega}_{j}(\mathbf{k})),$ (20)
where the primed summation is taken over all wavevectors in the first Brillouin zone and the sum over $j$ is taken over all surface unit cell phonon modes. Equations 19 and 20 simply separate the sum in Eqs. 15 and 16 into two parts: the outer sum varying intra-cell displacements and the inner sum varying inter-cell displacements. In principle, this procedure eliminates the dependence on boundary conditions and generalizes results from a surface unit cell to an infinite periodic surface. However, the depth of the surface (corresponding to the non-periodic dimension) is still limited to the depth of the surface unit cell.
Figure 4: Friction kernel for a model adsorbate on Pt(111) in a finite system (black) and in the infinite system size limit (blue). The finite system is equivalent to the 4x4x8 results presented in Figure 1.
Figure 4 demonstrates the results of Eqs. 19 and 20 and compares to previous results for a single unit cell slab. Naturally, the strong coupling of the adsorbate to a single acoustic mode is broadened, resulting in a much flatter spectral density in the low frequency ranges. Such a flat spectral density is characteristic of Markovian (white) noise. The high frequency range of the spectral density is largely unaltered, due to the dispersionless nature of the high frequency modes. In the SI, we discuss how a flat spectral density for acoustic phonons is consistent with predictions from continuum elastic theory.
4 Phonon effects on reaction rates
One of the primary motivations for studying the atomic vibrations of a solid surface is to elucidate the role of phonons in reactions rates at catalytic interfaces. As was noted first by Kramers44, and expanded on by many others45, 46, 47, 48, the friction of the environment plays a critical role in determining reaction rates, and especially in determining rate prefactors. In this section, we use transition state theory (TST) to demonstrate how surface phonons affect reaction rates, emphasizing the role of the phonon friction kernel. We then specialize to the case of desorption rates and compare the results of our model to experimental temperature dependent rate constants for CO and Xe desorption from Pt(111). Several details of the formalism given this section have been omitted for the purpose of providing a concise presentation of the theory that focuses on physical ramifications. A thorough presentation of the theory is provided in the SI for interested readers.
4.1 Transition State Theory
4.1.1 General theory
The Eyring transition state theory (TST) equation for the rate constant may be expressed as,
$k=\frac{1}{\beta h}\frac{Q^{{\ddagger}}}{Q_{R}}e^{-\beta\Delta E^{{\ddagger}}},$ (21)
where $h$ is Planck’s constant, $\Delta E^{{\ddagger}}$ is the potential energy difference between the reactant state and the transition state, $Q^{{\ddagger}}$ is the partition function at the transition state, and $Q_{R}$ is the partition function in the reactant basin. By expanding said partition functions to second order about the transition state and the reactant basin, respectively, we can generate analytical expressions for $Q_{R}$ and $Q^{{\ddagger}}$ yielding,
$k=\frac{1}{2\pi}\times\frac{\prod\limits^{N-1}_{i=0}f_{i}}{\prod\limits^{N-1}_{i=1}f_{i}^{{\ddagger}}}\times e^{-\beta\Delta E^{{\ddagger}}},$ (22)
where $N$ is the total number of modes, $f_{i}$ are the vibrational frequencies in the reactant basin, $f_{i}^{{\ddagger}}$ are the non-imaginary vibrational frequencies at the transition state. Note that since the transition state is defined to be a saddle point, it always has one imaginary frequency mode and this mode corresponds the reaction coordinate. Pollak illustrated that Eq. 22 is equivalent to Kramers and Grote-Hynes rate theories49. In a gas phase molecular reaction, $f_{i}$ is the $i$th eigenfrequency of the molecular mass-weighted Hessian. However, in a reaction at a surface interface, both surface phonon modes and molecular/adsorbate modes will contribute to the rate constant. We thus separate the product of eigenfrequencies into contributions from the solid surface and the molecular adsorbate to yield,
$k=\frac{1}{2\pi}\times\prod\limits^{N_{S}-1}_{i=0}\frac{\omega_{i}}{\omega_{i}^{{\ddagger}}}\times\frac{\prod\limits^{N_{A}-1}_{i=0}\tilde{f}_{i}}{\prod\limits^{N_{A}-1}_{i=1}\tilde{f}^{{\ddagger}}_{i}}\times e^{-\beta\Delta E^{{\ddagger}}},$ (23)
where $N_{S}$ and $N_{A}$ are the number of solid and adsorbate degrees of freedom respectively ($N_{S}+N_{A}=N)$), $\omega_{i}$ and $\omega_{i}^{{\ddagger}}$ are the phonon frequencies in the reactant and transition state respectively, and $\tilde{f}_{i}$ and $\tilde{f}^{{\ddagger}}_{i}$ are effective molecular frequencies in the reactant and transition state respectively. Let $\mathbf{H}_{A}$ be the mass-weighted Hessian of the adsorbate degrees of freedom, then $\tilde{f}_{i}$ and $\tilde{f}^{{\ddagger}}_{i}$ are the eigenfrequencies of an effective Hessian,
$\tilde{\mathbf{H}}_{A}=\mathbf{H}_{A}-\mathbf{C}\boldsymbol{\omega}^{-2}\mathbf{C}^{T},$ (24)
where the shift term is equal to the instantaneous ($t=0$) friction, $\mathbf{K}(t=0)=\mathbf{C}\boldsymbol{\omega}^{-2}\mathbf{C}^{T}$.
Comparing Eqns. 22 and 23 reveals that phonons introduce two distinct corrections to rate constant: the ratio of solid phonon frequencies between the reactant and transition state, and the shift in adsorbate frequencies. The ratio of phonon frequencies is largest when the reactant molecules are strongly coupled to the surface and the transition state is not. In Figure 2C we demonstrated that an adsorbate strongly coupled to a surface affects only the highest frequency phonon mode, leaving the bulk of the phonon density of states unchanged. Thus, we can approximate the ratio of phonon frequencies as,
$\prod\limits^{N_{S}-1}_{i=0}\frac{\omega_{i}}{\omega_{i}^{{\ddagger}}}\approx\frac{\tilde{\omega}_{\mathrm{D}}}{\omega_{\mathrm{D}}},$ (25)
where $\tilde{\omega}_{\mathrm{D}}$ is the highest frequency phonon mode when the reactants are bound to the surface and $\omega_{\mathrm{D}}$ is the bare solid Debye frequency.
The shift in the adsorbate frequencies is a thermodynamic correction arising from phonons altering the free energy surface along the reaction coordinate. Indeed, if we denote $\tilde{f}_{0}$ as the adsorbate normal mode along the reaction coordinate Eq. 23 can be simplified to,
$k=\frac{\tilde{f}_{0}}{2\pi}\times\prod\limits^{N_{S}-1}_{i=0}\frac{\omega_{i}}{\omega_{i}^{{\ddagger}}}\times e^{-\beta(\Delta E^{{\ddagger}}+T\Delta\tilde{S}^{\ddagger})},$ (26)
where $\Delta\tilde{S}^{\ddagger}$ is the effective barrier entropy,
$\Delta\tilde{S}^{\ddagger}=k_{\mathrm{B}}\sum^{N_{A}-1}_{i=0}\ln\left(\frac{\tilde{f}^{{\ddagger}}_{i}}{\tilde{f}_{i}}\right).$ (27)
In the harmonic approximation this barrier entropy is independent of temperature, but in a more general context it can be shown to be temperature dependent.
4.1.2 Surface desorption
The formalism described in the previous subsection can be applied to derive a rate constant for surface desorption. For gas-phase desorption, the reaction coordinate can be defined as the distance of the adsorbate center of mass from the surface. Furthermore, in many cases, the desorption process is barrierless, meaning that at the transition state the molecule and surface do not interact 50. Thus, we can adapt Eq. 25 to express the gas-phase desorption rate constant,
$k_{d}=\frac{\tilde{\omega}_{\mathrm{as}}}{2\pi}\times\frac{\tilde{\omega}_{\mathrm{D}}}{\omega_{\mathrm{D}}}\times e^{-\beta\Delta E^{{\ddagger}}},$ (28)
where $\tilde{\omega}_{\mathrm{as}}$ is the effective adsorbate-surface interaction frequency satisfying,
$\tilde{\omega}_{\mathrm{as}}=\sqrt{\frac{\mu}{m}\omega^{2}_{\mathrm{as}}-K(t=0)}.$ (29)
While transition state theory is primarily a classical theory 51, 52, 53, 54, studies have demonstrated improved agreement with experiment when introducing quantum corrections, such as accounting for the rotational motion of the molecule or using quantum harmonic oscillator partition functions instead of classical oscillator partition functions 50, 55. We will thus compare four different models for the desorption rate constant to experimental results: (1) a fixed-surface model using classical harmonic oscillator partition functions and a rotation correction,
$k_{d1}=\frac{\omega_{\mathrm{as}}}{2\pi}\times\frac{2I}{\hslash^{2}\beta}\times e^{-\beta\Delta E^{{\ddagger}}},$ (30)
where $I$ denotes the moment of inertia of the adsorbate. (2) a phonon-corrected model using classical harmonic oscillator partition functions and a rotational correction,
$k_{d2}=\frac{\tilde{\omega}_{\mathrm{as}}}{2\pi}\times\frac{\tilde{\omega}_{\mathrm{D}}}{\omega_{\mathrm{D}}}\times\frac{2I}{\hslash^{2}\beta}\times e^{-\beta\Delta E^{{\ddagger}}},$ (31)
(3) a fixed-surface model using quantum harmonic oscillator partition functions and a rotational correction,
$k_{d3}=\frac{1-e^{-\beta\hslash\omega_{\mathrm{as}}}}{2\pi\beta\hslash}\times\frac{2I}{\hslash^{2}\beta}\times e^{-\beta\Delta E^{{\ddagger}}},$ (32)
and (4) a phonon-corrected model using quantum harmonic oscillator partition functions and a rotational correction,
$k_{d4}=\frac{1-e^{-\beta\hslash\tilde{\omega}_{\mathrm{as}}}}{2\pi\beta\hslash}\times\frac{1-e^{-\beta\hslash\tilde{\omega}_{\mathrm{D}}}}{1-e^{-\beta\hslash\omega_{D}}}\times\frac{2I}{\hslash^{2}\beta}\times e^{-\beta\Delta E^{{\ddagger}}}.$ (33)
Note that by "fixed-surface" we do not mean a surface at absolute zero, but rather a surface that acts as an ideal, structureless, thermal environment.
4.2 Results and discussion
We have compared results from Eqs. 30 to  33 to experimental temperature-dependent desorption rate constants for CO and Xe from a Pt(111) surface. Parameters used in computing Eqs. 30 to  33 are presented in Table 1. Note that because Xe is an atomic species, its moment of inertia is 0, and thus we ignore the rotational partition function factors for Xe calculations. The phonon corrections in Eq. 31 and Eq. 33 were computed using a 4x4x8 EMT surface slab and subsequently averaged across the first Brillouin zone to approximate an infinite surface, as described in Sec. 3. In the SI, we illustrate that the rate constant corrections we present here are not sensitive to the size of the surface slab used, or whether one accounts for phonon dispersion.
Table 1: Parameters used for computing desorption rate constants shown in Figure 5.
$E^{{\ddagger}}$ ($\mathrm{eV}$)
$\omega_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
$\tilde{\omega}_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
$\tilde{\omega}_{\mathrm{D}}$ ($\mathrm{cm^{-1}}$)
CO
1.4755
48032
164
203
Xe
0.24556
2834
21
156
The CO desorption rate constants were taken from Ref. 55 and the Xe desorption rates were taken from Ref. 56. In Ref. 55, desorption rate constants were calculated by fitting the time-dependent flux from a beam scattering experiment to two models: a single exponential model and a bi-exponential model. The single exponential model fit the entire flux signal, mixing contributions from terrace and steps. Meanwhile the biexponential model separated the flux into a fast component, arising from terrace desorption, and a slow component, arising from step to terrace diffusion followed by terrace desorption. While our TST calculations do not include the role of steps, we compare the results of our models to data both from the single exponential model and the fast component of the biexponential for thoroughness and transparency.
Figure 5: Rate constants for desorption from a Pt(111) surface. (A) CO desorption. Grey squares are experimental data which mixed contributions from both steps and terraces. Black circles refer experimental data where the kinetics of terrace desorption was isolated. (B) Xe desorption.
In Fig. 5A, we see that the phonon-corrected models (Eqs. 31 and  33) give improved agreement with experimental results for CO desorption. In particular, $k_{d4}$ the quantum, flexible surface model and the terrace desorption rate constants yield the best agreement. The improved agreement when using Eqs. 31 to  33 versus Eqs. 30 to  32 arises from the reduced adsorbate-surface frequency $\tilde{\omega}_{\mathrm{as}}$. Physically, the flexible surface reduces the stiffness of the adsorbate-surface bond, leading to a lower frequency of attempts over the barrier and a lower rate prefactor.
In Fig. 5B, we demonstrate results for Xe. Here, all the TST models lie essentially on top of each other and are lower in value than the experimental rate constants, although by a small margin. The smaller phonon corrections for Xe versus CO are a natural result of the weaker interactions with the surface. Weaker coupling means a small $K(t=0)$, leading to $\tilde{\omega}_{\mathrm{as}}\approx\omega_{\mathrm{as}}$. Furthermore, weak coupling also results in a phonon density of states that is unchanged from a bare lattice, implying $\tilde{\omega}_{\mathrm{D}}=\omega_{\mathrm{D}}$. In general, the theory presented in this section suggests the phonon corrections to the rate constant are much larger for chemisorbed species than physisorbed species.
It is worth emphasizing that using a slightly different values for the surface binding energy, $\Delta E^{{\ddagger}}$, can substantially shift the quality of agreement of theoretical calculations with experiment. The major impediment to the first principles calculation of chemical rates is still the calculation of the barrier energy, and the phonon corrections to the rate constants seem to be a comparatively minor factor, even for chemisorbed species. Indeed, the purpose of this section was not to demonstrate that the magnitude of phonon corrections to reaction rates is large, but rather to illustrate that the theoretical models we developed in Sections 2 and 3, when combined with transition state theory, produce physically interpretable results which correspond well with existing experimental measurements.
5 Methods
All calculations were performed using the Atomic Simulation Environment (ASE) python package 57, 58, 59, 60, 61, 62. Mass-weighted Hessians were computed using a finite-difference scheme in the vibrations module of ASE. Dynamical matrices were computed using the phonon module. The displacement size used in finite difference calculations was 0.01. 100 points were used to uniformly sample the first Brillouin zone.
When computing the friction kernels and spectral densities, all adsorbates were treated as an effective adatom, with a given coupling between the center of mass of the adsorbate and a surface atom. Friction kernels were computed for the three Cartesian degrees of freedom of this adatom. Additionally, a single atom in the bottom layer of each surface slab was constrained in order to remove center of mass motion. Without a constraint on the solid, the center of mass diffusion of the entire surface will dominate the contribution to the friction kernel. Results for friction kernels were computed using individual surface atoms as adsorption sites and subsequently averaged.
6 Conclusions
In this manuscript we have developed a theory for how surface phonons couple to molecular adsorbates based on the generalized Langevin equation. By integrating out the solid degrees of freedom (assuming they could be described harmonically) we derived a GLE for the adsorbate, wherein the friction is merely a sum of the phonon frequency of the solid weighted by their coupling to the adsorbate. We demonstrated that this friction kernel depends sensitively on the frequency of the adsorbate-surface bond. When the frequency of this bond is smaller than the Debye frequency of the solid, adsorbates couple primarily to the acoustic phonons of solid. When the frequency of the bond is much larger than the Debye frequency of the solid, adsorbates couple primarily to the dispersionless local vibrations of the adsorption site. Subsequently, we used harmonic transition state theory to derive phononic corrections to reaction rate constants. We show that these corrections improved agreement between theory and experiment for CO desorption rates from Pt(111).
Supplementary Material
See supplementary material for results for friction kernels for surfaces other than Pt(111), a perturbative analysis of the physisorbed and chemisorbed limits, more details on convergence of the friction kernel with respect to supercell size, a discussion of continuum elastic theory, and more details on the derivation of rate constants using transition state theory.
Data Availability
[…]
Acknowledgements
AF and APW were supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-SC0019441. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Ardavan Farahvash acknowledges support from the National Science Foundation Graduate Research Fellowship program.

### v3

@@
 In this manuscript, we provide a general theory for how surface phonons couple to molecular adsorbates.
-Our theory maps the extended dynamics of a solid’s atomic vibrational motions to a generalized Langevin equation, and by doing so captures these dynamics in a single quantity: the non-Markovian friction.
-The different frequency components of this friction are the phonon modes of the solid weighted by their coupling to the adsorbate degrees of freedom.
+Our theory maps the extended dynamics of a surface’s atomic vibrational motions to a generalized Langevin equation, and by doing so captures these dynamics in a single quantity: the non-Markovian friction.
+The different frequency components of this friction are the phonon modes of the surface slab weighted by their coupling to the adsorbate degrees of freedom.
 Using this formalism, we demonstrate that physisorbed species couple primarily to acoustic phonons while chemisorbed species couple to dispersionless local vibrations.
@@
 The central parameter of this equation is the non-Markovian friction kernel (or memory kernel), which encapsulates the different surface vibrational modes weighted by their coupling to the adsorbate.
+This friction kernel can be developed to represent surface modes with wavelengths well beyond that of a typical simulation cell, thereby eliminating a common source of finite-size effects.
 Using this formalism, we demonstrate that the influence of surface phonons on the dynamics of an adsorbed molecule depends significantly on the magnitude of the adsorbate-surface coupling.
 Specifically, we find that chemisorbed species couple primarily to dispersionless local vibrations, while physisorbed species couple primarily to acoustic phonons across a broad range of frequencies.
-The key parameter that determines whether the primary influence of surface vibrations is via extended or localized phonon modes is the ratio of the solid’s Debye frequency to the frequency of the adsorbate-surface bond (Figure 1).
-By combining these observations with harmonic transition state theory, we derive equations that describe how phonons alter reaction rates at solid surfaces and demonstrate that these phononic corrections agree with experimental measurements of desorption rate constants.
+The key parameter that determines whether the primary influence of surface vibrations is via extended or localized phonon modes is the ratio of the solid’s Debye frequency to the frequency of the adsorbate-surface bond (Fig.
+1).
+By combining these observations with harmonic transition state theory, we derive equations that describe how phonons alter reaction rates at solid surfaces, and demonstrate that these phononic corrections agree with experimental measurements of desorption rate constants.
+The approach we present in this manuscript extends a previous method for utilizing the GLE for describing the influence of surface vibrational modes on molecular adsorbates.
+27, 28 In the previous method, the GLE was designed to describe the collective influences of the surrounding solid on a single representative surface atom (i.e., binding site).
+In the method we present here, the GLE is designed to describe the influence of the entire surface on the dynamics of the adsorbate itself.
+This new approach is more easily interpretable because the friction kernel directly represents influence of surface phonons on the adsorbate.
 Figure 1: 2D schematic illustrating the dominant phonon modes in terms of their coupling to the adsorbate or contribution to the friction kernel.
 $\omega_{\mathrm{as}}$ is the frequency of the adsorbate-surface bond and the $\omega_{\mathrm{D}}$ is the solid’s Debye frequency.
-The approach we present in this manuscript extends our previous method for utilizing the GLE for describing the influence of surface vibrational modes on molecular adsorbates.
-farahvash_using_2023 In our previous method, the GLE was designed to describe the collective influences of the surrounding solid on a single representative surface atom (i.e., binding site).
-In the method we present here, the GLE is designed to describe the influence of the entire surface on the dynamics of the adsorbate itself.
-This new approach is more easily interpretable because the friction kernel directly represents influence of surface phonons on the adsorbate.
 The remainder of the paper is organized as follows.
@@
 2.1 Theory
-In this section, we derive a theory for modeling the effects of solid surface vibrations on the dynamics of an adsorbate.
+In this section, we derive a theory for modeling the effects of surface vibrations on the dynamics of an adsorbate.
 The general approach involves integrating out the nuclear degrees of freedom of the solid and representing their influence as effective equations of motion for the adsorbate.
@@
 First, we assume that the nuclear degrees of freedom of the adsorbate and solid can be described classically and adiabatically.
-There are cases where adsorbate dynamics are non-adiabatic, and in such cases a mixed quantum-classical GLE approach has been shown to be effective 27, 28, however the development and analysis of such a theory is beyond the scope of this paper.
+There are cases where adsorbate dynamics are non-adiabatic, and in such cases a mixed quantum-classical GLE approach has been shown to be effective 29, 30, however the development and analysis of such a theory is beyond the scope of this paper.
 Second, we assume that the solid potential energy surface and surface-adsorbate interaction are both harmonic.
-This approximation simplifies the formulation and is appropriate for the development of generalizable physical insight.
+This approximation simplifies the formulation of the theory and is appropriate for the development of generalizable physical insight.
 We note that the effects of anharmonicity can be incorporated into the theory, which would be a natural target of future developments.
@@
 8 and Eq.
-9 yields a GLE for the adsorbate degrees of freedom29, 30,
+9 yields a GLE for the adsorbate degrees of freedom31, 32,
 […]
@@
 For example, consider the case of an adsorbate that interacts with a single surface atom via a single adsorbate atom (e.g.
-CO on Pt(111) 31, 32).
+CO on Pt(111) 33, 34).
 If the force constant associated with this interaction is $\mu\omega_{\mathrm{as}}^{2}$, where $\mu$ is the reduced mass of the adsorbate and surface atom, Eqs.
@@
 We have computed the phonon friction kernel for a simple model of CO on Pt(111).
-The CO was modeled as a single adatom and assumed the interact with a single adsorption site, as experimental structures show that CO adsorbs primarily atop Pt(111) sites 31, 32.
+The CO was modeled as a single adatom and assumed the interact with a single adsorption site, as experimental structures show that CO adsorbs primarily atop Pt(111) sites 33, 34.
 We artificially vary the adsorbate-surface frequency $\omega_{\mathrm{as}}$ over a range of physically motivated values, in order asses how the strength of the adsorbate-surface interaction affects the properties of the friction kernel.
-The metal potential energy surface $V_{S}$ was modeled using an effective medium theory (EMT) forcefield developed by Norskov et al.33 More details about calculations can be found in the Section 5.
+The metal potential energy surface $V_{S}$ was modeled using an effective medium theory (EMT) forcefield developed by Norskov et al.35 More details about calculations can be found in the Section 5.
 Figure 2: (A) Friction kernel, (B) spectral density, and (C) density of states for Pt(111) and three values of $\omega_{\mathrm{as}}$.
 The density of states in (C) was calculated using the 4x4x8 surface slab.
-Figure 2 illustrates the results of our calculations for a selected set of $\omega_{\mathrm{as}}$ values ranging from those characteristic of weakly adsorbed species to those characteristic of strongly adsorbed species.
+Fig.
+2 illustrates the results of our calculations for a selected set of $\omega_{\mathrm{as}}$ values ranging from those characteristic of weakly adsorbed species to those characteristic of strongly adsorbed species.
 We plot the friction kernel computed in the direction normal to the surface in systems with either a 4x4x8 or 8x8x8 Pt(111) slab.
 The lower values of $\omega_{\mathrm{as}}$ that we consider lie below that of the platinum’s Debye frequency of $\omega_{\mathrm{D}}=156\mathrm{cm}^{-1}$.
-We associate these values with the physisorbed regime, as they are characteristic of weakly adsorbed systems such as noble gases on Pt(111).34 The higher values of $\omega_{\mathrm{as}}$ that we consider are associated with the chemisorbed regime, corresponding to physical systems such as CO on Pt(111), which has $\omega_{\mathrm{as}}\approx 480\ \mathrm{cm}^{-1}$).32, 35, 36 Delta function peaks in $\bar{K}(\omega)$ and $\rho(\omega)$ were broadened to thin Lorentzians of width $1\mathrm{cm}^{-1}$ for ease of visualization.
+We associate these values with the physisorbed regime, as they are characteristic of weakly adsorbed systems such as noble gases on Pt(111).36 The higher values of $\omega_{\mathrm{as}}$ that we consider are associated with the chemisorbed regime, corresponding to physical systems such as CO on Pt(111), which has $\omega_{\mathrm{as}}\approx 480\ \mathrm{cm}^{-1}$).34, 37, 38 Delta function peaks in $\bar{K}(\omega)$ and $\rho(\omega)$ were broadened to thin Lorentzians of width $1\mathrm{cm}^{-1}$ for ease of visualization.
+It is important to note that the non-decaying oscillations in $K(t)$ are a consequence of the harmonic approximation.
+Accounting for anharmonicities will cause the memory kernel to asymptotically decay at long-times.
 For all cases of $\omega_{\mathrm{as}}$, the adsorbate couples strongly to a surface acoustic mode appearing around $\omega=10\mathrm{cm}^{-1}$.
@@
 The phonon mode associated with this peak is associated with the flexing of the lattice in the direction perpendicular to the surface.
-Snapshots of this motion are provided in the insets in Figure 1 and video animations are provided in the electronic SI.
+Snapshots of this motion are provided in the insets in Fig.
+1 and video animations are provided in the electronic SI.
 In the chemisorbed regime, the friction kernel is dominated by a high frequency mode around $\omega=210\ \mathrm{cm}^{-1}$.
@@
 The large amplitude of this mode signifies that in the chemisorbed limit, the adsorbate is primarily sensitive to the local oscillations of the surface binding site (which are shifted in frequency due to the presence of the adsorbate).
-Many published models for reactive scattering on solid surfaces will describe solid vibrations using only a single, harmonically bound surface atom 6, 37, 38, 39, 40.
+Many published models for reactive scattering on solid surfaces will describe solid vibrations using only a single, harmonically bound surface atom 6, 39, 40, 41, 42.
 These results explain why such a method is successful for strongly coupled species.
-Despite this large shift in the spectral density when varying $\omega_{\mathrm{as}}$, in Figure 2C we demonstrate that the phonon density of states is nearly identical to that of a bare surface.
+Despite this large shift in the spectral density when varying $\omega_{\mathrm{as}}$, in Fig 2C we demonstrate that the phonon density of states is nearly identical to that of a bare surface.
 The only significant change in $\rho(\omega)$ is the presence of the aforementioned surface site local mode present around $\omega=210\ \mathrm{cm}^{-1}$.
@@
 In the SI, we rigorously examine this statement by comparing perturbative schemes to the exact results for $K(t)$ and $\bar{K}(\omega)$.
-Finally, we note that in this section we have only studied the effects of phonons for an ideal clean surface.
+Perturbation theory also allows us to derive analytical approximations to the friction kernel (Eqs.
+S10-S12) and precisely define the "phase boundaries" in Figure 1.
+We acknowledge that in this section we have only studied the effects of phonons for an ideal clean surface.
 Presumably, sources of surface heterogeneity, such as steps and defects, may add further richness and complexity to the picture we have provided here.
-We also note that we have only studied elemental solids, and the optical modes of polyatomic crystals could add another interesting dimension to this physical picture.
+We also note that we have only studied elemental solids, and the optical modes of polyatomic crystals could add another interesting dimension to our physical picture.
+Finally, in this section we have only studied species at a gas-solid interface.
+Many important surface-chemical processes occur at liquid-solid interfaces, where it has been demonstrated sorption dynamics play a critical role in electrokinetic transport and measurements 43.
 3 Phonon dispersion
@@
 In a surface slab, the presence of anisotropy breaks the symmetry of the 3D crystal.
-This symmetry breaking leads to additional surface mode bands, which can be acoustic (Rayleigh waves), or non-acoustic in character41, 42, 43.
+This symmetry breaking leads to additional surface mode bands, which can be acoustic (Rayleigh waves), or non-acoustic in character44, 45, 46.
 In Fig.
@@
 We verified convergence with respect to supercell size.
-Of course, the number of surface phonon bands depends on the size of the surface unit cell used in computing $\mathbf{D}(\mathbf{k})$, however, we demonstrate in in Figure S7 results are qualitatively similar across different unit cell sizes.
+Of course, the number of surface phonon bands depends on the size of the surface unit cell used in computing $\mathbf{D}(\mathbf{k})$, however, we demonstrate in in Fig.
+S7 results are qualitatively similar across different unit cell sizes.
 Figure 3: Phonon dispersion curves.
@@
 Figure 4: Friction kernel for a model adsorbate on Pt(111) in a finite system (black) and in the infinite system size limit (blue).
-The finite system is equivalent to the 4x4x8 results presented in Figure 1.
+The finite system is equivalent to the 4x4x8 results presented in Fig.
+2.
 Figure 4 demonstrates the results of Eqs.
@@
 One of the primary motivations for studying the atomic vibrations of a solid surface is to elucidate the role of phonons in reactions rates at catalytic interfaces.
-As was noted first by Kramers44, and expanded on by many others45, 46, 47, 48, the friction of the environment plays a critical role in determining reaction rates, and especially in determining rate prefactors.
+As was noted first by Kramers47, and expanded on by many others48, 49, 50, 51, the friction of the environment plays a critical role in determining reaction rates, and especially in determining rate prefactors.
 In this section, we use transition state theory (TST) to demonstrate how surface phonons affect reaction rates, emphasizing the role of the phonon friction kernel.
@@
 Note that since the transition state is defined to be a saddle point, it always has one imaginary frequency mode and this mode corresponds the reaction coordinate.
-Pollak illustrated that Eq.
-22 is equivalent to Kramers and Grote-Hynes rate theories49.
 In a gas phase molecular reaction, $f_{i}$ is the $i$th eigenfrequency of the molecular mass-weighted Hessian.
@@
 where the shift term is equal to the instantaneous ($t=0$) friction, $\mathbf{K}(t=0)=\mathbf{C}\boldsymbol{\omega}^{-2}\mathbf{C}^{T}$.
-Comparing Eqns.
+Comparing Eqs.
 22 and 23 reveals that phonons introduce two distinct corrections to rate constant: the ratio of solid phonon frequencies between the reactant and transition state, and the shift in adsorbate frequencies.
 The ratio of phonon frequencies is largest when the reactant molecules are strongly coupled to the surface and the transition state is not.
-In Figure 2C we demonstrated that an adsorbate strongly coupled to a surface affects only the highest frequency phonon mode, leaving the bulk of the phonon density of states unchanged.
+In Fig.
+2C we demonstrated that an adsorbate strongly coupled to a surface affects only the highest frequency phonon mode, leaving the bulk of the phonon density of states unchanged.
 Thus, we can approximate the ratio of phonon frequencies as,
@@
 where $\tilde{\omega}_{\mathrm{D}}$ is the highest frequency phonon mode when the reactants are bound to the surface and $\omega_{\mathrm{D}}$ is the bare solid Debye frequency.
+The factor of $\tilde{\omega}_{\mathrm{D}}$ is directly related to our observation that chemisorbed species couple strongly to local modes above the natural Debye frequency of the metal.
+The outsized influence of these modes on the dynamics of the chemisorbed species reflects why it appears in the expression for the rate constant.
+For physisorbed species $\frac{\tilde{\omega}_{\mathrm{D}}}{\omega_{\mathrm{D}}}=1$.
 The shift in the adsorbate frequencies is a thermodynamic correction arising from phonons altering the free energy surface along the reaction coordinate.
@@
 where $\Delta\tilde{S}^{\ddagger}$ is the effective barrier entropy,
-$\Delta\tilde{S}^{\ddagger}=k_{\mathrm{B}}\sum^{N_{A}-1}_{i=0}\ln\left(\frac{\tilde{f}^{{\ddagger}}_{i}}{\tilde{f}_{i}}\right).$ (27)
+$\Delta\tilde{S}^{\ddagger}=k_{\mathrm{B}}\sum^{N_{A}-1}_{i=10}\ln\left(\frac{\tilde{f}^{{\ddagger}}_{i}}{\tilde{f}_{i}}\right).$ (27)
 In the harmonic approximation this barrier entropy is independent of temperature, but in a more general context it can be shown to be temperature dependent.
+In Eqs.
+23 and 26, the influence of the phonon friction kernel on the rate constant is both explicit and implicit.
+It is explicit in terms of $K(t=0)$, but also implicit in ratio of normal mode frequencies.
+The dependence of this ratio on the friction kernel can only be made explicit if additional simplifications are made.
+Notably, under the condition that the coupling between the adsorbate degrees of freedom and the solid degrees of freedom $\mathbf{C}$ is the same in both the reactant basin and transition state, harmonic TST reduces to Kramers-Grote-Hynes(KGH) theory, where the rate prefactor is explicitly related to the Laplace transform of the friction kernel 52.
+However, such a condition is not appropriate even in simple desorption processes, and therefore we do not directly use KGH theory here.
+See the SI (Section S6) for additional details on KGH theory.
 4.1.2 Surface desorption
@@
 For gas-phase desorption, the reaction coordinate can be defined as the distance of the adsorbate center of mass from the surface.
-Furthermore, in many cases, the desorption process is barrierless, meaning that at the transition state the molecule and surface do not interact 50.
+Furthermore, in many cases, the desorption process is barrierless, meaning that at the transition state the molecule and surface do not interact 53.
 Thus, we can adapt Eq.
@@
 $\tilde{\omega}_{\mathrm{as}}=\sqrt{\frac{\mu}{m}\omega^{2}_{\mathrm{as}}-K(t=0)}.$ (29)
-While transition state theory is primarily a classical theory 51, 52, 53, 54, studies have demonstrated improved agreement with experiment when introducing quantum corrections, such as accounting for the rotational motion of the molecule or using quantum harmonic oscillator partition functions instead of classical oscillator partition functions 50, 55.
+While transition state theory is primarily a classical theory 54, 55, 56, 57, studies have demonstrated improved agreement with experiment when introducing simple quantum corrections, such as accounting for the rotational motion of the molecule or using quantum harmonic oscillator partition functions instead of classical oscillator partition functions 53, 58.
 We will thus compare four different models for the desorption rate constant to experimental results: (1) a fixed-surface model using classical harmonic oscillator partition functions and a rotation correction,
@@
 30 to  33 are presented in Table 1.
-Note that because Xe is an atomic species, its moment of inertia is 0, and thus we ignore the rotational partition function factors for Xe calculations.
+We ignore the rotational partition function factors for Xe calculations.
 The phonon corrections in Eq.
@@
 In the SI, we illustrate that the rate constant corrections we present here are not sensitive to the size of the surface slab used, or whether one accounts for phonon dispersion.
-Table 1: Parameters used for computing desorption rate constants shown in Figure 5.
-$E^{{\ddagger}}$ ($\mathrm{eV}$)
-$\omega_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
-$\tilde{\omega}_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$)
-$\tilde{\omega}_{\mathrm{D}}$ ($\mathrm{cm^{-1}}$)
-CO
-1.4755
-48032
-164
-203
-Xe
-0.24556
-2834
-21
-156
+Table 1: Parameters used for computing desorption rate constants shown in Fig. 5.
+$E^{{\ddagger}}$ ($\mathrm{eV}$) $\omega_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$) $\tilde{\omega}_{\mathrm{as}}$ ($\mathrm{cm^{-1}}$) $\tilde{\omega}_{\mathrm{D}}$ ($\mathrm{cm^{-1}}$)
+CO 1.4758 48034 164 203
+Xe 0.24559 2836 21 156
 The CO desorption rate constants were taken from Ref.
 58 and the Xe desorption rates were taken from Ref.
-56.
+59.
 In Ref.
-55, desorption rate constants were calculated by fitting the time-dependent flux from a beam scattering experiment to two models: a single exponential model and a bi-exponential model.
+58, desorption rate constants were calculated by fitting the time-dependent flux from a beam scattering experiment to two models: a single exponential model and a bi-exponential model.
 The single exponential model fit the entire flux signal, mixing contributions from terrace and steps.
@@
 In general, the theory presented in this section suggests the phonon corrections to the rate constant are much larger for chemisorbed species than physisorbed species.
+We also note that the deviation between theory and experiment in Fig.
+5B is larger for lower temperatures.
+For sufficiently low-mass species at sufficiently cryogenic temperatures, it has been shown that tunneling plays a significant role in barrier crossing60, requiring more sophisticated quantum corrections to the rate constants than those in Eq.
+32 and Eq.
+33.
 It is worth emphasizing that using a slightly different values for the surface binding energy, $\Delta E^{{\ddagger}}$, can substantially shift the quality of agreement of theoretical calculations with experiment.
@@
 5 Methods
-All calculations were performed using the Atomic Simulation Environment (ASE) python package 57, 58, 59, 60, 61, 62.
+All calculations were performed using the Atomic Simulation Environment (ASE) python package 61, 62, 63, 64, 65, 66.
 Mass-weighted Hessians were computed using a finite-difference scheme in the vibrations module of ASE.